OpenStudy (anonymous):

Not really math sorry, still love the help! Suppose 128 radioactive atoms have a half-life of 15 seconds. Approximately how many radioactive atoms will be left after 60 seconds? A. 32 B. 16 C. 8 D. None I was told you can do this after 15, 64 will be left after 30, 32 will be left after 45, 16 after 60, 8 8 is it I was wondering if I could just divide 128 / 15. I did and got 8.

OpenStudy (anonymous):

no you cannot do 128/15. that wouldn't give you 8 in any case. but your answer is right.

OpenStudy (anonymous):

Haha damn it. >.<

OpenStudy (anonymous):

half life means that in a certain amount of time, half the original number of atoms would decay. a half life of 15 seconds means that the original number of atoms would decay to half the original number every 15 seconds. so, starting with 128 atoms, after 15 seconds, you have 64 atoms after another 15 seconds you have 32 atoms after another 15 seconds you have 16 atoms after another 15 seconds you have 8 atoms. the correct formula then is n = 60/15 = 4 number of atoms after 1 minute = 128/2^n = 128/(2^4) = 128/16 = 8